OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1) _ _ _ y = _

OpenStudy (neo92):

This can be done in two ways. 1). let the eqn of the line be \[y = mx + c\] substitute the given co-ordinates and solve the two simultaneous eqns for m an c 2). use the fact that gradiend of a st.line is constant \[(x-x^{'})/(y-y^{'}) = (x^{''}-x^{'})/(y^{''}-y^{'})\] where x',x'',y',y'' are the cordinates of the given points

OpenStudy (neo92):

got it?

OpenStudy (anonymous):

What would M be?

OpenStudy (neo92):

what did u get?

OpenStudy (anonymous):

Nevermind.

OpenStudy (neo92):

didnt u try? should i give u some time o continue?

OpenStudy (neo92):

ok. the answer is /[7y + x = 12/]